You need 400,000 to buy a flat. So, you go to the bank and ask for that amount. The bank offers an annual interest rate of 3%, and the loan has to be paid back in 20 years, in constant monthly payments. Calculate how much you'll have to pay back to the bank every month (write the result with no decimal points) Answer:
How much you should invest today in a financial product that offers a 5% interest rate (compounded yearly) in order to be able to afford six yearly payments, the first one of them of 25,000 EUR to be done next year, and further payments increasing in a 3% per year? (write the result with no decimal points) Answer:
You own an oil pipeline that will generate a 2M cash return over the coming year. The pipeline's operating costs are negligible, and it is expected to last forever. Unfortunately, the volume of oil shipped is declining, and cash flow are expected to decline by 4% per year. The discount rate is 10%. What is the PV of the pipeline's cash flows? (write the result with no decimal points) Answer
l shall buy some shares that will produce cash-flows (incomes) of 10,000 for life. In six years I will earn a first amount of 10,000. The profitability for this type of investments is 5%. Calculate how much should i pay today for these shares (write the result with no decimal points) Answer:
Answer #1
1) Amount to be paid back to the bank every month = 400000*0.0025*1.0025^240/(1.0025^240-1) = $              2,218
2) The amount to be deposited = 25000/1.05+25000*1.03/1.05^2+25000*1.03^2/1.05^3+25000*1.03^3/1.05^4+25000*1.03^4/1.05^5+25000*1.03^5/1.05^6 = $        1,36,225
3) PV of the pipeline's cash flow = 2000000/(0.10+0.04) = $ 1,42,85,714
4) The amount to be paid = 10000/0.05+10000/1.05^6 = $        2,07,462